# MgFNO: Multi-grid Architecture Fourier Neural Operator for Parametric Partial Differential Equations

### v2

@@ Abstract
In science and engineering, there is often a need to repeatedly solve large-scale and high-resolution partial differential equations (PDEs). Neural operators are a new model that can map between function spaces, allowing trained models to emulate PDE solution operators. This paper introduces a novel Fourier neural operator with a multigrid architecture (MgFNO). The MgFNO combines the frequency principle of deep neural networks (DNNs) with the multigrid idea of solving linear systems. To speed up the FNO training process, a three-layer V-cycle multigrid architecture is used. This architecture involves training the model multiple times on a coarse grid and then transferring it to a fine grid to accelerate the training of the model. The solver based on DNN learns the solution from low to high frequency, while the multigrid method acquires the solution from high to low frequency. Note that the FNO is a resolution-invariant solution operator, so the corresponding calculations are greatly simplified. Finally, experiments are carried out on Burgers’ equation, Darcy flow, and Navier-Stokes equation. The results show that the proposed MgFNO outperforms the traditional Fourier neural operator. Code Availability: https://github.com/guozihao-hub/MgFNO/tree/master.
@@ 1 Introduction
In practice, this linear system is typically large-scale and is often solved using iterative methods. Commonly employed iteration methods include splitting iterative methods such as the Jacobi iteration, Gaussian-Seidel iteration, and successive over-relaxation (SOR) iteration, as well as Krylov subspace methods such as the conjugate gradient method (CG) and generalized minimal residual (GMRES) method [4].
@@
Traditional numerical methods for PDEs solve equations in discrete space. Consequently, the traditional numerical solution of partial differential equations entails a trade-off in the selection of the discrete resolution. In other words, a coarse grid (low resolution) has fast calculation speed but low accuracy, whereas a fine grid (high resolution) has high accuracy but slow calculation speed. In practice, for complex partial differential equation systems, very fine discretization is usually required, which requires a huge amount of time and computational resources, making it difficult to solve partial differential equations [5].
@@
The multigrid method (MG) [6, 7] is one of the most remarkable classical methods. The computational complexity of the MG method for Poisson’s equation is only $O(n)$, where $n$ is the matrix size. Another popular high-performance numerical method is the spectral method, which has a complexity of $O(nlogn)$.
@@
Recently, with the explosive development of computing resources and data sampling technology, machine learning and data-driven methods, especially deep neural networks, have rapidly developed in the field of information data and have been widely applied to image recognition and natural language processing problems [8, 9]. However, when analyzing complex physical, biological, or engineering systems, the cost of data collection is often prohibitively high, and we inevitably face the challenge of making decisions based on partial information. In small-data systems, the vast majority of state-of-the-art machine learning techniques (such as deep/convolutional/recursive neural networks) lack robustness and cannot guarantee convergence. To address this issue, Physical Information Neural Networks (PINNS) [10] have been proposed, which directly parameterize the solution function into a neural network. Palha and Rebelo [10] point out that for many physics related modeling problems, there is often a large amount of prior knowledge that has not yet been used in modern machine learning practices. This prior knowledge can play the role of regularization constraints, constraining the space of tolerable solutions to a manageable range. As a reward, encoding this structured information into the learning algorithm amplifies the data seen by the algorithm, enabling it to quickly converge to the correct solution, and even in situations where only a small amount of training data can be obtained, the machine algorithm still maintains good generalization ability [11].
@@
Machine learning methods can provide fast solvers that approximate or enhance traditional disciplines, thereby becoming the key to the scientific discipline revolution. However, PINNS maps between finite dimensional spaces; therefore, they can only learn solutions related to specific discretization. This is usually a limitation in practical applications. Therefore, it is necessary to develop grid-independent neural networks to solve partial differential equation systems. Meanwhile, for a new instance of a differential system, this method requires retraining the model and is limited to the setting of known potential PDEs. In response to these issues, Lu and Pang [12] proposed the operator neural network (DeepONet), which compensates for the mesh dependency of the finite-dimensional operator method by generating a single network parameter set that can be used with different discretizations. It can transfer solutions between grids with different resolutions. In addition, the operator neural networks need to be trained only once. The solution for obtaining new parameter instances only requires forward propagation of the network, thereby alleviating the main computational problems that arise in neural finite element methods. Finally, operator neural networks do not require knowledge of underlying PDEs, only data, and are completely data-driven methods. Due to the high cost of calculating integral operators, DeepONet has not yet produced effective numerical algorithms, and therefore cannot achieve successful replication of convolutional or recursive neural networks in finite-dimensional environments. In another approach, a graph neural operator (GNO) was proposed [13]. GNO focuses on learning infinite-dimensional mapping by the composition of nonlinear activation functions and a certain class of integral operators. Subsequently, Li et al. [14] proposed a Fourier neural operator (FNO) based on this foundation. FNO uses fast Fourier transform (FFT) to approximate partial differential equations, which is more in line with the laws of physical equations than convolutional and recursive neural networks, and has shown good performance and speed in flow problems such as Darcy flow and N-S equations. Zhang and Zhao [15] has developed a deep learning based FNO model to solve three types of problems in PDE control of two-dimensional oil/water two-phase flow underground. His work has shown great potential in replacing traditional numerical methods with neural networks for reservoir numerical simulations.
@@
When constructing a network model, the FNO needs to specify the parameter ”modes,” which represents the number of frequency components retained during the Fourier transform. If only a small number of low-frequency components are retained, the training time of the neural network model is short, but the training error is often large. If more frequency components are retained, the training error is relatively small, but the model training time is often longer. Figure 1 illustrates the trade-off between computation speed and fitting accuracy imposed by FNO. Another shortcoming of the FNO is that the basis functions in the Fourier transform are generally frequency-localized. This implies that a specific basis function corresponds to a specific frequency, resulting in clear and concentrated basis functions in the frequency domain. While the Fourier transform is effective for analyzing the frequency components of signals, it is not suitable for representing the local transformations of signals in time or space. This limitation has led to the development of other transformations, such as the wavelet transform. The wavelet neural operator (WNO) [16] is proposed to overcome this aforementioned shortcoming. It transforms snapshots into high- and low-frequency components using discrete wavelet transform.
@@
In this paper, our goal is to enhance the training process of FNO by focusing on frequency aspects. To begin with, we propose a three-layer V-cycle multigrid architecture for training the model. This involves training the model multiple times on a coarse grid and then transferring it to a fine grid to expedite the training process. The use of multigrid effectively addresses the challenge of eliminating high-frequency components. The underlying concept is that high-frequency components on fine grids can be transformed into low-frequency components on coarse grids. Existing research has already explored the integration of multigrid methods to develop efficient DNNs. For example, in [20], a unified model called MgNet is introduced, which combines convolutional neural networks (CNN) for image classification with multigrid (MG) methods for solving discretized partial differential equations (PDEs). Another study [21] proposes a linear U-net structure as a solver for linear PDEs on a regular mesh. Additionally, we incorporate compact support activation functions to ensure that FNO achieves rapid and consistent convergence across multiple scales [22]. Using the aforementioned observations and concepts, we developed a novel Fourier neural operator known as MgFNO. This innovative approach addresses the issue of slow training speed in FNO from a frequency perspective. Moreover, our proposed MgFNO has proven to outperform the traditional Fourier neural operator in solving various equations, including the 1-D Burgers’ equation, 2-D Darcy flow, and 2-D time-dependent Navier-Stokes equations.
The remainder of this paper is organized as follows. In Section 2, we introduce the necessary notations for the rest of the paper and review the use of the MG method and FNO to solve parametric PDEs. In Section 3, we discuss the frequency principle and various methods for overcoming low-frequency trends. We focus on the construction of a multiscale DNN that can handle the challenge of converging high-frequency components for neural networks. In Section 4, the proposed MgFNO method is introduced. Section 5 presents numerical experiments and comparisons among the classical FNO methods. Finally, concluding remarks are presented in section 6.
@@ 2 The MG method and FNO to solve parametric PDEs
In this section, we briefly describe the multigrid method and Fourier neural operator to solve the PDEs. MG method is one of the most efficient methods for solving the PDEs. FNO is a new and effective framework for solving PDEs, which surpasses existing neural network methods in benchmark PDE problems.
@@ 2.1 The Multigrid method
Before using the MG method to solve the PDEs (1.1), it is necessary to discretize the PDEs. The finite difference method (FDM), which directly approximates differential equations on spatial grids, is one of the most commonly used discretization methods. Using differential discretization, each node is approximated, and the differential operator is replaced by a differential approximation. In the finite element method (FEM), the solution area is first divided into a finite number of small elements, and appropriate mathematical interpolation functions are used within each element to approximate the solution. By introducing shape functions and element stiffness matrices, the behavior of the overall system can be transformed into relationships between local elements. The finite volume method (FVM) focuses on conserving fluxes by integrating governing equations over discrete control volumes. This approach ensures that the flux entering and leaving the control volume is accurately accounted for, thereby preserving the overall balance in the system. In summary, both methods discretize a continuous problem into a discrete problem and then obtain the numerical solution of the problem through numerical methods.
For a linear system (1.2), the iterative method is a basic numerical method. Let the initial guess value of the solution be $u_{0}$ and the update scheme be represented by $\mathscr{H}$. We can write the iterative method
@@
Figure 2 illustrates the convergence factor distribution of Jacobi’s method for solving linear systems of the Poisson equation. It is evident that iterative methods (2.1) effectively reduce high-frequency errors; however, they are generally ineffective at addressing low-frequency errors. This is due to the localized nature of iterative methods, which struggle to eliminate low-frequency global error components that arise from the overall pattern.
To overcome the limitations of iterative methods (2.1), R. P. Fedorenko in the 1960s proposed the MG method and further developed by Achi Brandt and others in the 1970s. The core idea of this method is to use grids of different scales to accelerate convergence. By solving the problem on a coarse grid, low-frequency errors can be eliminated more quickly, whereas a fine grid is used to improve the accuracy of the solution and eliminate high-frequency errors. The MG method can be divided into two steps: fine-grid smoothing and coarse-grid correction.
The fine-mesh smoothing process uses a smoother to eliminate high-frequency errors. The smoother is usually an iterative method, which can effectively eliminate high-frequency approximation errors. However, it is not very effective in eliminating low-frequency errors for which coarse grid corrections are required. The MG method approximates low-frequency errors using solutions on a coarse grid. The computational complexity of the MG method is typically $O(N)$, where $N$ denotes the number of grid points. This is much more efficient than the traditional iterative methods that have complexities of $O(N^{2})$ or $O(N^{3})$. The specific MG algorithm is presented in Algorithm 1.
Figure 3: : Comparisons of computational complexity of different methods.
Algorithm 1 MG method
@@ 3 Frequency Tendency in Deep Learning and Compensation Methods
Although deep neural networks (DNNs) possess powerful capabilities, their theoretical analysis is not yet sufficient; therefore, DNNs are also known as ”black boxes.” However, in recent years, some studies have revealed this magical ”black box” by demonstrating the frequency principle of deep neural network training behavior. In this section, we first introduce low-frequency bias in the neural network fitting process. Subsequently, we explore the utilization of multiscale neural networks and the conversion of coarse and fine grids as approaches to overcome the high-frequency curse.
@@ 3.1 Frequency principle
Recently, several papers [17, 18, 19] have discovered that DNNs exhibit frequency-dependent convergence behavior. Specifically, during training, DNNs tend to fit target functions from low to high frequencies. This phenomenon is known as the F-Principle of DNNs[17]. The slow convergence of DNNs in the later stages of training may be attributed to the challenges in learning high-frequency components of the data. In order to demonstrate the F-Principle, we utilize 1-D synthetic data to observe the evolution of a DNN in both the frequency and spatial domains. Our objective is to train a DNN to fit a 1-D target function $f(x)=sin(x)+sin(5x)$, which comprises two frequency components. The data set consists of 1,000 evenly distributed sampling points within the interval $[-2\pi,2\pi]$. We compute the discrete Fourier transform of $f(x)$ using the formula $\hat{f}_{k}=\frac{1}{n}\textstyle\sum_{i=0}^{n-1}f(x_{i})e^{-j2\pi ik/n}$, where $k$ represents the frequency. We initialize the parameters of the DNN using a Gaussian distribution with a mean of 0 and a standard deviation of 0.1. We employ a ReLU-DNN with widths of 1-200-200-1, and set the learning rate to 0.0005. Figure 5 and Figure 6 illustrate the rapid convergence of the DNN for low frequencies, while the convergence is significantly slower for high frequencies.
Figure 5: : Training process of DNNs in the spatial domain. Training data are sampled from the target function $f(x)=sin(x)+sin(5x)$, see [19].
@@
It is worth noting that the iterative method (2.1) to solve linear systems often converges first on the high-frequency component of the error and then on the low-frequency component. In other words, the convergence tendency of iterative methods is exactly opposite to that of DNNs. The MG method improves upon the iterative method by utilizing the characteristic that the low-frequency components on the coarse grid are high-frequency components on the fine grid. This has inspired us, and we have also attempted to use this feature to construct a multigrid model to accelerate the fitting process of FNO.
Next, we will define a neural network and explain the concept of the frequency principle. First, let $T(x)=Wx+b$ be a linear function, where $W=(w_{ij})\in\mathbb{R}^{m\times n}$ represents the weights, and $b\in\mathbb{R}^{m}$ represents the biases. Then, let $\sigma(u):\mathbb{R}\rightarrow\mathbb{R}$ be a nonlinear activation function. Common activation functions include the sigmoid, Tanh, and ReLU functions. Finally, a DNN $NN(x)$ with $L$ layers can be expressed as
@@
This theorem reveals the relative rate of change between the low-frequency and high-frequency components of the loss function during training. It indicates that when the gradient descent method is applied to the loss function, the low-frequency part of the loss function converges faster than the high-frequency part.

### v3

@@ Abstract
Neural operators are a new type of models that can map between function spaces, allowing trained models to emulate the solution operators of partial differential equations(PDEs). This paper proposes a multigrid Fourier neural operator (MgFNO) that accelerates the training of traditional Fourier neural operators through a novel three-level hierarchical architecture. The key innovation of MgFNO lies in its decoupled training strategy employing three distinct networks at different resolution levels: a coarse-level network first learns low-resolution approximations, an intermediate network refines the solution, and a fine-level network achieves high-resolution accuracy. By combining the frequency principle of deep neural networks with multigrid methodology, MgFNO effectively bridges the complementary learning patterns of neural networks (low-to-high frequency) and multigrid methods (high-to-low frequency error reduction).Experimental results demonstrate that MgFNO achieves relative errors of 0.17%, 0.28%, and 0.22% on the Burgers’ equation, Darcy flow, and Navier-Stokes equations, respectively, representing reductions of 89%, 71%, and 83% compared to the conventional FNO. Furthermore, MgFNO supports zero-shot super-resolution prediction, enabling direct application to high-resolution scenarios after training on coarse grids. This study establishes an efficient and high-accuracy new paradigm for solving complex PDEs dominated by high-frequency dynamics. Code and data used are available on https://github.com/guozihao-hub/MgFNO/tree/master.
@@ 1 Introduction
This linear system is usually large-scale in practice and often solved using iterative methods. Commonly employed iteration methods include splitting iterative methods such as Jacobi iteration, Gaussian-Seidel iteration, and successive over-relaxation (SOR) iteration, as well as Krylov subspace methods such as the conjugate gradient method (CG) and the generalized minimal residual (GMRES) method [4].
@@
Traditional numerical methods for PDEs solve equations through discrete space. Consequently, the traditional numerical solution of partial differential equations entails a trade-off in the selection of discrete resolution. That is to say, a coarse grid (low resolution) has a fast calculation speed but low accuracy, whereas a fine grid (high resolution) has high accuracy but slow calculation speed. In practice, for complex partial differential equation systems, very fine discretization is usually required, which requires a huge amount of time consumption and computational resources, which brings practical difficulties to the solution of partial differential equations [5].
@@
The multigrid method (MG) [6, 7] is one of the most remarkable classical methods. The computational complexity of the MG method for Poisson’s equation is only $O(n)$, where $n$ is the size of the matrix. Another popular high-performance numerical method is the spectral method, which has a complexity of $O(nlogn)$.
@@
Recently, with the explosive development of computing resources and data sampling technology, machine learning and data-driven methods, especially deep neural networks, have rapidly developed in the field of information data and have been widely applied to image recognition and natural language processing problems [8, 9]. However, when analyzing complex physical, biological, or engineering systems, the cost of data collection is often prohibitively high, and we inevitably face the challenge of making decisions based on partial information. In small data systems, the vast majority of state-of-the-art machine learning techniques (such as deep/convolutional/recursive neural networks) lack robustness and cannot provide any guarantee of convergence. To address this issue, Physical Information Neural Networks (PINNS) [10] is proposed, which directly parameterizes the solution function into a neural network. Palha and Rebelo [10] point out that for many physics related modeling problems, there is often a large amount of prior knowledge that has not yet been used in modern machine learning practices. These prior knowledge can play the role of regularization constraints, constraining the space of tolerable solutions to a manageable range. As a reward, encoding this structured information into the learning algorithm amplifies the data seen by the algorithm, enabling it to quickly converge to the correct solution, and even in situations where only a small amount of training data can be obtained, the machine algorithm still maintains good generalization ability [11].
@@
Machine learning methods can provide fast solvers that approximate or enhance traditional disciplines, thereby becoming the key to the scientific discipline revolution. However, PINNS maps between finite dimensional spaces. For a new instance of a differential system, this method requires retraining the model and is limited to the setting of known potential PDEs. In response to the above issues, Lu and pang [12] proposed the Operator neural network (DeepONet), which compensates for the mesh dependency of the finite dimensional operator method by generating a single network parameter set that can be used with different discretizations. It can transfer solutions between grids of different resolutions. In addition, operator neural networks only need to be trained once. The solution for obtaining new instances of parameters only requires forward propagation of the network, thereby alleviating the main computational problems that arise in neural finite element methods. Finally, operator neural networks do not require knowledge of underlying PDEs, only data, and are completely data-driven methods. Due to the high cost of calculating integral operators, DeepONet has not yet produced effective numerical algorithms, and therefore cannot achieve successful replication of convolutional or recursive neural networks in finite-dimensional environments. In a different approach the graph neural operator (GNO) was proposed [13]. The GNO focuses on learning of the infinite-dimensional mapping by composition of nonlinear activation functions and certain class of integral operators. Subsequently, Li, Kovachki and Azizzadenesheli [14] proposed the fourier neural operator (FNO) based on this foundation. FNO uses fast fourier transform (FFT) to approximate partial differential equations, which is more in line with the laws of physical equations than convolutional and recursive neural networks, and has shown good performance and speed in flow problems such as Darcy flow and N-S equations. Zhang and Zhao [15] has developed a deep learning based FNO model to solve three types of problems in PDE control of two-dimensional oil/water two-phase flow underground. His work has shown great potential in replacing traditional numerical methods with neural networks for reservoir numerical simulation.
@@
When constructing a network model, FNO needs to specify the parameter ”modes”, which represents the number of frequency components retained during the Fourier transform. If only a small number of low-frequency components are retained, the training time of the neural network model is short, but the training error is often large. If more frequency components are retained, the training error is relatively small, but the model training time is often longer. Figure 1 illustrates the trade-off between computation speed and fitting accuracy imposed by FNO. Another shortcoming of FNO is that the basis functions in the Fourier transform are generally frequency localized. This means that a specific basis function corresponds to a specific frequency, resulting in clear and concentrated basis functions in the frequency domain. While the Fourier transform is effective at analyzing the frequency components of signals, it is not suitable for representing local transformations of signals in time or space. This limitation has led to the development of other transformations, such as the wavelet transform. The wavelet neural operator (WNO) [16] is proposed to overcome this aforementioned shortcoming. It transforms the snapshots into their high and low frequency components using discrete wavelet transform.
@@
In this paper, our goal is to enhance the training process of FNO by focusing on frequency aspects. To begin with, we propose a three-layer V-cycle multigrid architecture for training the model. This involves training the model multiple times on a coarse grid and then transferring it to a fine grid to expedite the training process. The use of multigrid effectively addresses the challenge of eliminating high-frequency components. The underlying concept is that high-frequency components on fine grids can be transformed into low-frequency components on coarse grids. Existing research has already explored the integration of multigrid methods to develop efficient DNNs. For example, in [20], a unified model called MgNet is introduced, which combines convolutional neural networks (CNN) for image classification with multigrid (MG) methods for solving discretized partial differential equations (PDEs). Another study [21] proposes a linear U-net structure as a solver for linear PDEs on a regular mesh. Additionally, we incorporate compact support activation functions to ensure that FNO achieves rapid and consistent convergence across multiple scales [22]. Using the aforementioned observations and concepts, we have developed a novel Fourier neural operator known as MgFNO. This innovative approach addresses the issue of slow training speed in FNO from a frequency perspective. Moreover, our proposed MgFNO has proven to outperform the traditional Fourier neural operator in solving various equations, including the 1-D Burgers’ equation, 2-D Darcy flow, and 2-D time-dependent Navier-Stokes equation.
The remaining sections are organized as follows. In section 2, we introduce necessary notations for the rest of the paper and review the use of the MG method and FNO to solve parametric PDEs. In section 3, we discuss the frequency principle and various methods to overcome low-frequency trends. We pay special attention to the construction of a multiscale DNN that can handle the challenge of converging high-frequency components for neural networks. In section 4, we introduce the proposed MgFNO. Section 5 presents the numerical experiments and comparisons among the classical FNO methods. Finally, concluding remarks are given in section 6.
@@ 2 The MG method and FNO to solve parametric PDEs
In this section, we briefly describe multigrid method and fourier neural operator to solve the PDEs. MG method is one of the most efficiency methods for solving PDEs. FNO is a new and effective framework for solving PDEs, which surpasses the existing neural network methods in the benchmark PDE problems.
@@ 2.1 The Multigrid method
Before using the MG method to solve PDEs (1.1), it is necessary to discretize the PDEs. The finite difference method (FDM)is one of the most commonly used discretization methods, which directly approximates differential equations on spatial grids. By using differential discretization, each node is approximated and the differential operator is replaced by differential approximation. In the finite element method (FEM), the solution area is first divided into a finite number of small elements, and appropriate mathematical interpolation functions are used within each element to approximate the solution. By introducing shape functions and element stiffness matrices, the behavior of the overall system can be transformed into relationships between local elements. The finite volume method (FVM) focuses on conserving fluxes by integrating the governing equations over discrete control volumes. This approach ensures that the flux entering and leaving a control volume is accurately accounted for, preserving the overall balance in the system. In summary, both of these methods discretize a continuous problem into a discrete problem, and then obtain the numerical solution of the problem through numerical methods.
For linear system (1.2), iterative method is one of the basic numerical methods. Let the initial guess value of the solution be $u_{0}$, and update scheme is represented by $\mathscr{H}$. We can write the iterative method as
@@
Figure 2 illustrates the distribution of convergence factors for the Jacobi iterative method when solving linear systems derived from the Poisson equation. It can be clearly observed that the iterative method (2.1) effectively reduces high-frequency errors but generally performs poorly in handling low-frequency errors. This limitation arises from the localized nature of iterative methods, which struggle to eliminate low-frequency global error components generated by overall patterns. Notably, the frequency characteristics of iterative methods exhibit opposite convergence tendencies compared to Fourier neural operators. Therefore, we aim to explore improvements to iterative methods, thereby enhancing the performance of Fourier neural operators.
To address the limitations of iterative methods (2.1), the multigrid (MG) method was initially proposed by R. P. Fedorenko in the 1960s and subsequently refined by Achi Brandt and others in the 1970s. The fundamental principle of this approach lies in utilizing hierarchical grid structures to accelerate convergence through two key algorithmic steps: (1) fine grid smoothing and (2) coarse grid correction. In the fine grid smoothing phase, conventional iterative methods (smoothers) are employed to effectively eliminate high-frequency approximation errors. However, recognizing that such local operations remain ineffective for low-frequency error components, the method strategically employs coarse grid correction - where low-frequency errors are approximated and corrected through solutions computed on coarser grids. This dual-scale mechanism enables the MG method to simultaneously achieve both rapid error reduction through coarse grid processing and high precision through fine grid refinement, with the complete algorithmic procedure detailed in Algorithm 1.
Algorithm 1 MG method[6]
@@ 3 Frequency Tendency in Deep Learning and Compensation Methods
Although deep neural networks (DNNs) possess powerful capabilities, their theoretical analysis is not yet sufficient, so DNNs are also known as ”black boxes”. However, in recent years, some studies have revealed this magical ”black box” by demonstrating the frequency principle of deep neural network training behavior. In this section, we first introduced the low-frequency bias in the neural network fitting process. Subsequently, we will explore the utilization of multiscale neural networks and the conversion of coarse and fine grids as approaches to overcome the high-frequency curse.
@@ 3.1 Frequency principle
Recently, several papers[17, 18, 19] have discovered that DNNs exhibit frequency-dependent convergence behavior. Specifically, during training, DNNs tend to fit target functions from low to high frequencies. This phenomenon is known as the F-Principle of DNNs[17]. The slow convergence of DNNs in the later stages of training may be attributed to the challenges in learning high-frequency components of the data. In order to demonstrate the F-Principle, we utilize 1-D synthetic data to observe the evolution of a DNN in both the frequency and spatial domains. Our objective is to train a DNN to fit a 1-D target function $f(x)=sinx+sin5x$, which comprises two frequency components. The data set consists of 1,000 evenly distributed sampling points within the interval $[-2\pi,2\pi]$. We compute the discrete Fourier transform of $f(x)$ using the formula $\hat{f}_{k}=\frac{1}{n}\textstyle\sum_{i=0}^{n-1}f(x_{i})e^{-j2\pi ik/n}$, where $k$ represents the frequency. We initialize the parameters of the DNN using a Gaussian distribution with a mean of 0 and a standard deviation of 0.1. We employ a ReLU-DNN with widths of 1-200-200-1, and set the learning rate to 0.0005. Figure 4 and Figure 5 illustrate the rapid convergence of the DNN for low frequencies, while the convergence is significantly slower for high frequencies.
Figure 4: : Training process of DNNs in the spatial domain.
@@
It is worth noting that the iterative method (2.1) to solve linear systems often converges first on the high-frequency component of the error, and then on the low-frequency component. In other words, the convergence tendency of iterative methods is exactly opposite to that of DNNs. The MG method improves upon the iterative method by utilizing the characteristic that the low-frequency components on the coarse grid are the high-frequency components on the fine grid. This has inspired us, and we have also attempted to use this feature to construct a multigrid model to accelerate the fitting process of FNO.
Next, we will define a neural network and explain the concept of the frequency principle. Firstly, let $T(x)=Wx+b$ be a linear function, where $W=(w_{ij})\in\mathbb{R}^{m\times n}$ represents the weights and $b\in\mathbb{R}^{m}$ represents the biases. Then, let $\sigma(u):\mathbb{R}\rightarrow\mathbb{R}$ be a nonlinear activation function. Common activation functions include the sigmoid function, Tanh function, ReLU function, etc. Finally, a DNN $NN(x)$ with $L$ layers can be expressed as:
@@
This theorem reveals the relative rate of change between the low-frequency and high-frequency components of the loss function during training. This theorem indicates that when the gradient descent method is applied to the loss function, the low-frequency part of the loss function converges faster than the high-frequency part.